OpenStudy (anonymous):

Which of the following statements best describes the graph of 4x - y = 1?

OpenStudy (anonymous):

It is a straight line joining the points (0, -1), (1, 3), and (-1, -5). It is a straight line joining the points (4, -3), (-1, 2), and (-4, 5). It is a curve joining the points (4, -1), (2, 3), and (4, 1). It is a curve joining the points (0, -1), (-1, -3), and (1, 5).

OpenStudy (anonymous):

@Hero @Luigi0210 @iambatman @kropot72

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

To see if any point lies on 4x - y = 1, you just plug in the coordinates and see if you get a true equation

jimthompson5910 (jim_thompson5910):

for example, let's check (x,y) = (0,-1) 4x - y = 1 4(0) - (-1) = 1 ... plug in x = 0 and y = -1 0 + 1 = 1 1 = 1 so (0,-1) lies on this line

OpenStudy (anonymous):

ok thnks

jimthompson5910 (jim_thompson5910):

so can you tell me if (1,3) lies on 4x - y = 1 or not?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

its 1,1 as well

OpenStudy (anonymous):

Answer is A @jim_thompson5910 ?

jimthompson5910 (jim_thompson5910):

you are correct

jimthompson5910 (jim_thompson5910):

it's a straight line

OpenStudy (anonymous):

Can you help me with 1 more?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

Sharon wants to make a graph to show the relationship between the number of tickets sold and the price of tickets sold. She plots the following points. (0, 0), (1, 20), (2, 40), (3, 60), (4, 80) She uses the following steps to plot the graph: Label origin as (0, 0) On x-axis, label Number of tickets sold. On y-axis, label Price of tickets in dollars. The scale on x-axis starts from 0 and goes up to 10 at intervals of 2. The scale on the y-axis starts from 0 and goes up to 200 at intervals of 20. Which of the following best describes the graph: It will not be spread out across the entire coordinate plane because in Step 1, Sharon selected an incorrect point as the origin. It will not be spread out across the entire coordinate plane because in Step 2, Sharon plotted the dependent variable on the x-axis. It will not be spread out across the entire coordinate plane because in Step 3, Sharon plotted the independent variable on the y-axis. It will not be spread out across the entire coordinate plane because in Step 4, and Step 5, Sharon selected incorrect scales on the axes.

jimthompson5910 (jim_thompson5910):

did you graph this?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

alright go ahead and do so

OpenStudy (anonymous):

ok i got it

jimthompson5910 (jim_thompson5910):

so based on what you have, what do you think the answer is and why?